OpenStudy (anonymous):

Need help with an statistical question! Two dependable probability variables X and Y, who have the standard deviation of o-x=0.87 o-y=0.38 and the correlation coefficient equals 0.09. Calculate the standard deviation of X+Y. The answer should be 0.9802

OpenStudy (anonymous):

By "o-x" and "o-y" I assume you mean \(\sigma_X\) and \(\sigma_Y\), respectively, as in the standard deviations of \(X\) and \(Y\)?

OpenStudy (anonymous):

yes excactly!

OpenStudy (anonymous):

For two dependent random variables \(X\) and \(Y\), the variance of their sum is equal to the sum of their variances as well twice the covariance: \[V(X+Y)=V(X)+V(Y)+2\text{ Cov}(X,Y)\] where the covariance can be expressed in terms of the the given correlation coefficient: \[\text{Cov}(X,Y)=\rho_{X,Y}\sigma_X\sigma_Y\] (\(\rho_{X,Y}\) denotes the correlation coefficient.) On top of that, the standard deviation is equivalent to square root of the variance. So, you have \[\begin{align*}V(X+Y)&=V(X)+V(Y)+2\rho_{X,Y}\sigma_X\sigma_Y\\\\ {\sigma_{X+Y}}^2&={\sigma_X}^2+{\sigma_Y}^2+2\rho_{X,Y}\sigma_X\sigma_Y\\\\ \sigma_{X+Y}&=\sqrt{(0.87)^2+(0.38)^2+2(0.09)(0.87)(0.38)}\\\\ &=0.9802 \end{align*}\]

OpenStudy (anonymous):

Thanks you so much for the explantation!